OpenStudy (anonymous):

For f(x)=1/x^2-3, find f(3), and f(2+h)

OpenStudy (anonymous):

Given: f(x)=1/x^2-3 Find: f(2+h) then, f(2+h)=1/(2+h)^2-3

OpenStudy (anonymous):

f(3)= 1/3^2-3 = 1/9-3 =1/6 f(2+h) = 1/[(2+h)^2-3] = 1/(4+4h+h^2-3) = 1/(1+4h+h^2)

OpenStudy (anonymous):

So f(3) = 1/6? what about the second one?

OpenStudy (anonymous):

showed you the second one : f(2+h) = 1/[(2+h)^2-3] = 1/(4+4h+h^2-3) = 1/(1+4h+h^2)